OpenStudy (anonymous):

use the distributive property to simplify 5(1/30+1/80)

OpenStudy (fibonaccichick666):

so, can you tell me what the distributive property is?

OpenStudy (anonymous):

Distributive property is (a+b)*c=ac+bc 5(1/30+1/80)=5/30+5/80=1/6+1/16

OpenStudy (anonymous):

Find the LCM for 6 and 16 and then symplify